If you randomly select a card from a well-shuffled standard deck of 52 cards, what is the probability that the card you select is a 7 or 3? (Your answer must be in the form of a reduced fraction.)



1
Expert's answer
2022-05-18T06:42:30-0400

There are four 7s and four 3s in a deck.


"P(7s\cup 3s)=\dfrac{4+4}{52}=\dfrac{2}{13}"